CHIRAL RECOGNITION IN NEUTRAL AND IONIC MOLECULAR COMPLEXES

\underline{ANANYA SEN}, AUDE BOUCHET, VALERIA LEPERE, KATIA LE BARBU-DEBUS, and ANNE ZEHNACKER-RENTIEN, CNRS, Institut des Sciences Moléculaires d’Orsay (ISMO), UMR 8214, Orsay F-91405, and Univ. Paris-Sud, Orsay F-91405, France.

The behaviour of chiral molecules in the gas phase have been studied. The two pseudoenantiomers of Quinine, have been put intact in the gas phase using laser ablation and characterized using different spectroscopic techniques such as REMPI, LIF, IR-UV double resonance. VCD measurements have been performed in solution. From this comparative study one can conclude that the pseudoenantiomers behave differently both in gas phase and in solution^{a}.

aAnanya Sen, Aude Bouchet, Valeria Lepère, Katia Le Barbu-Debus, D. Scuderi, F. Piuzzi, and A. Zehnacker-Rentien, J. Phys. Chem. A, 2012, 116 (32), pp 8334-8344